A bullet is fired horizontally at 500 m/s at 1.5 m

Show work, please.
Answer #1

u = 500 m/s

h = 1.5 m

x = u*t

h = 1/2gt^2

t = sqrt(2h/g)

t = 0.55 s

x = u*t

x = 276.6 m